OpenStudy (anonymous):

h'(t)=? h(t)=2cot^2(πt +2)

OpenStudy (anonymous):

have u tryed it yet?

OpenStudy (anonymous):

Yeah... to be honest my brain isn't working right now

OpenStudy (anonymous):

Take a break then :-D

OpenStudy (anonymous):

You'll want to remember the chain rule here... for: \[u^n \] then: \[(n*u^{n-1})*\frac{d}{du}u\] Do the power rule, then take the derivative of what's inside the "shell"/"layer"/"whatever"

OpenStudy (anonymous):

@ agentx5 I know the chain rule, but like... okay.. \[h(t)=2\cot ^{2}(\pi t +2)\] \[h'(t)=4\cot(\pi t +2)*\cot(\pi t +2)*-\sec^2(\pi t +2)(\pi)\] ... is there like double chain rule? ... I'ma go sleep after this

OpenStudy (anonymous):

@agentx5 didn't tag properly

OpenStudy (anonymous):

.. -sec^2 is supposed to be -csc^2.